OpenStudy (anonymous):

find the equation in standard form which passes through (-2, 7) and (1,1)

OpenStudy (anonymous):

can someone please attempt at helping me or solving this porblem

OpenStudy (anonymous):

y-7/(7-1)=X+2/(-2-1)

OpenStudy (anonymous):

(y-7)/(7-1)=(X+2)/(-2-1)

OpenStudy (amistre64):

what is their definition of a "standard form"?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

ax+by=c

OpenStudy (amistre64):

standard form is defined as ax+by=c ; but alot of times they use the terminology to mean slope intercept form ....

OpenStudy (amistre64):

so its good to know which one they are refering to :)

OpenStudy (anonymous):

2X+Y=3

OpenStudy (anonymous):

yes slopeintercept form is what i think they are reffering to

OpenStudy (amistre64):

subtract your points and stack the value of y over x like this: y/x to determine the slope (-2, 7) -(1, 1) ------ -3, 6 ; y/x = 6/-3, simplify to -2 right?

OpenStudy (amistre64):

now we can use a modified form developed from the point slope form of a line like this: y = mx -mPx +Py where m=slope; P(x,y) is either given point

OpenStudy (amistre64):

y = mx -mPx +Py ; (1,1) looks simple enough to use y = -2x --2(1) + 1 y = -2x +2 + 1 y = -2x +3

OpenStudy (anonymous):

finding m=(1-7)/(1--2) =-6/3 =-2 use: y=mx +c we find c=11 when sub:(-2;7) then y=2x+11

OpenStudy (amistre64):

we can test this one: y=2x+11 ; at (1,1) 1 = 2(1) + 11 1 = 2 + 11 1 = 13 .... aint it

OpenStudy (amistre64):

chk it with mine we get: y = -2x +3 ; (1,1) 1 = -2(1) + 3 1 = -2 + 3 1 = 1 ............. good y = -2x +3 ; (-2,7) 7 = -2(-2) + 3 7 = 4 + 3 7 = 7 ........... good

OpenStudy (anonymous):

y=.5x+.5

OpenStudy (amistre64):

mz ... care to prove that result?

OpenStudy (amistre64):

teen, do you see why my results are correct?

OpenStudy (anonymous):

THANK YOU ALL.. i have been following along with your work and have came up with an answer I'm confident with.

OpenStudy (anonymous):

y=mx+c m=(7-1)/(-2-1)=-2 y=-2x+c puttng 1,1 y=-2x+3 opppssss!!!!!

OpenStudy (amistre64):

lol .... at least you proofed it ;)

OpenStudy (anonymous):

sry@64